OpenStudy (mayaal):

Giving a medal and a fan! using the given zero,find all the other zeroes of f(x). -2i is a zero of f(x)=x^4-45x^2-196

OpenStudy (anonymous):

use inverse operations and distribute property and solve

OpenStudy (mayaal):

How?

OpenStudy (anonymous):

well use algebra if its easier for you and you use that because it makes the problem more easier to solve

OpenStudy (mayaal):

but what can i use to fin all the other zeroes?

OpenStudy (anonymous):

use the zero property and solve

OpenStudy (anonymous):

there are a couple of ways to do this problem

OpenStudy (anonymous):

ok

OpenStudy (mayaal):

OK,so for that,i will have to factor it,right?

OpenStudy (anonymous):

one way is that since \(-2i\) is a zero then so is \(2i\) that means that it must have a factor of \[(x+2i)(x-2i)=x^2+4\] so factor it as \[x^4-45x^2-196=(x^2+4)(something)\] and then find the zeros of the something using either the quadratic formula or by factoring it

OpenStudy (mayaal):

(x^2+4)(x^2-49),right?

OpenStudy (anonymous):

the other way is to view \(x^4-45x^2-196\) as a quadratic in \(x^2\)

OpenStudy (anonymous):

yeah that does it

OpenStudy (mayaal):

(x+7)(x-7)(x-2i)(x+2i)

OpenStudy (anonymous):

yup

OpenStudy (mayaal):

zeroes:-7,7,2i,-2i?

OpenStudy (anonymous):

yes

OpenStudy (mayaal):

Thnku very much @satellite73 :)

OpenStudy (anonymous):

yw you did most of the work

OpenStudy (mayaal):

:)

OpenStudy (mayaal):

and u 2 @herua123 :)

OpenStudy (anonymous):

:)